Find the product:
  (3/11) × (2/5)
Find the product:
  (3/11) × (2/5)
Find the product:
  (3/11) × (2/5)

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

The product of two rational numbers = (product of their numerator)/ (product of their denominator)
We have,
= (3×2)/ (11×5)
= 6/55

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve it as follows :
3/11 × 2/5
= (3×2)/(11×5)
= 6/55
 
Hope it helps!
Thank you ✌️